# Recursive Gaussian Process State Space Model

### v3

@@ 4.3 Algorithm Summary
In summary, we have addressed the four challenges outlined in the Introduction for implementing online GPSSMs. Specifically, first, nonlinearity and coupling are handled through first-order linearization and joint inference of the system state and GP, leading to a closed-form Bayesian update equation. Second, a dynamic inducing-point adjustment algorithm is developed to overcome the computational burden caused by the non-parametric nature of the GP. Third, we recover the likelihood model from the approximate posterior distribution and utilize it for the online optimization of GP hyperparameters, thus enhancing learning accuracy. Based on these solutions, we have developed an efficient online GPSSM method that can adapt to both the operating domain and GP hyperparameters. Due to its flexible recursive inference capability for GPSSMs, this method is referred to as the recursive GPSSM (RGPSSM), which is summarized in Algorithm 1. A schematic diagram of the algorithmic workflow is shown in Fig. 1.
@@
0:Input: initial inducing points $\bm{u}$, threshold for adding points $\varepsilon_{tol}$, budget for the inducing-point set $M$.
1:repeat
2:  Assess the novelty $\gamma$ of the new point $f_{t}$ using (23).
3:  if $\gamma>\varepsilon_{tol}$ then
4:   Add $f_{t}$ to the inducing-point set and propagate the moments using (17).
5:  else
6:   Propagate the moments using (24).
7:  end if
@@
10:  end if
11:  if a new measurement $y_{t}$ is available then
12:   Correct the moments using (18).
13:  end if
14:  Minimize $\mathcal{L}$ in (31) to update the GP hyperparameters $\bm{\theta}$ using the Adam optimizer [53], and subsequently update the moments using (32).
15:until operation ends
@@
In theory, Algorithm 1 has the recursive learning capability for GPSSMs; however, as with the Kalman filter, there are numerical stability issues when implemented on a computer. Specifically, the algorithm requires the propagation of the covariance matrix $\bm{\Sigma}_{t}$ of the augmented state $\bm{X}_{t}$, which can easily lose positive definiteness due to the accumulation of rounding errors. To address this problem, a stable implementation based on Cholesky factorization is derived, as detailed in Appendix 7.4. This implementation method is capable of enhancing the numerical stability by making the computation more compact. In the next section, we will discuss the computational complexity of the Cholesky version of RGPSSM, along with the extension to the multi-output GP cases.
@@ 5.1 Multi-output GP Version
where ${K}_{0}(\cdot,\cdot)$ in the RHS represents the single-output kernel, and $\bm{\Xi}$ denotes the signal covariance between each function output dimension. Based on this multi-output kernel function, the kernel matrix has a Kronecker product structure, i.e., $\bm{K}_{0}(\mathcal{X}_{1},\mathcal{X}_{2})={K}_{0}(\mathcal{X}_{1},\mathcal{X}_{2})\otimes\bm{\Xi}$. This structure indicates that the $(i,j)$-th $d_{f}\times d_{f}$ block of the kernel matrix $\bm{K}_{0}(\mathcal{X}_{1},\mathcal{X}_{2})$ is the prior covariance between function values $\bm{f}(\textbf{x}_{i})$ and $\bm{f}(\textbf{x}_{j})$. Therefore, the only modification from the single-output GP version of RGPSSM to the multi-output GP version is to convert scalar elements in related vectors or matrices into corresponding blocks, which are presented in the following.
@@
  • $\bm{\Omega}_{dd}$ is the $i_{\mathrm{d}}$-th diagonal block of the inducing points submatrix in $\Omega$, where the inducing points submatrix is defined as the matrix slice $\bm{\Omega}[d_{x}:,d_{x}:]$
@@
Overall, the extension to the multi-output GP version is straightforward, with modifications involving the kernel function, the definitions of some vectors and matrices, and the evaluation criteria for inducing points management. It is evident that the dimensions of statistical moments will scale with the dimension of GP output, thus impacting the computational complexity, which is analyzed in the next subsection.
@@ 5.2 Computational Complexity Analysis
The computational complexity of RGPSSM is analyzed across its various algorithmic steps in the following table, where a detailed comparison between the standard and Cholesky-based implementations of RGPSSM is presented. The analysis is based on the multi-output GP version of RGPSSM.
@@
Based on the above analysis, additional remarks regarding the computational complexity are presented below:
  • By precomputing and storing the Cholesky factor of the kernel matrix $\bm{K}_{uu}$ during updates, we can avoid repeated matrix inversions in GP prediction (10), thereby reducing the computational complexity to quadratic order.
@@
  • It is clear that the overall algorithmic complexity grows cubically with both the function output dimension $d_{f}$ and the number of inducing points $M$. This is mainly due to two factors: the EKF-like moment matching, which provides second-order learning capabilities [54], and the flexible inducing point adjustment and hyperparameter optimization algorithms. Although this cubic complexity limits the algorithm’s applicability to high-dimensional scenarios, it offers a stable, fast, and flexible learning solution for low-dimensional problems, which are common in many applications such as aerospace [8] and robotics [50].
@@ 5.3 Generalization and Applicability
RGPSSM can be regarded as a unified generalization of several existing methods. Specifically, it subsumes both the pre-parameterized AKF-based online GPSSM with inducing points [41] and kernel-based online regression approaches such as KRLS [51] and SOGP [52] as special cases. For example, if RGPSSM is initialized with a fixed set of inducing points and no further updates are performed (i.e., by setting $\varepsilon_{\text{tol}}=+\infty$), it reduces to the pre-parameterized AKF-based method. Furthermore, when the transition and measurement functions in the GPSSM are set as $x_{t}=f(\bm{c}_{t})$ and $y_{t}=x_{t}$, where $\bm{c}_{t}$ and $y_{t}$ denote the function input and output, respectively, RGPSSM becomes equivalent to KRLS and SOGP. Therefore, RGPSSM offers practical and flexible recursive learning capabilities applicable to a wide range of scenarios, as summarized below:
  • Applicable to several observation scenarios, including partial state observation ($\bm{y}_{t}=\bm{g}(\bm{x}_{t})$), full state observation ($\bm{y}_{t}=\bm{x}_{t}$), and direct function value observation ($\bm{y}_{t}=\bm{f}(\bm{x}_{t})$).
@@
  • The use of first-order linearization, together with the augmentation of the inducing point set, enables a finite-dimensional Gaussian approximation of the joint state–inducing point distribution. This provides a flexible foundation for incorporating advanced nonlinear filters, such as the sigma point filter [55, 56], iterative filter [57], variational filter [58], or exact moment evaluation [59], to further improve the accuracy of moment matching.
@@ 6 Experiment Result
This section demonstrates the proposed method using two synthetic simulations and a real-world dataset. Section 6.1 presents a comparative experiment on a synthetic NASCAR® dataset, which serves as a benchmark for online GPSSM methods. Section 6.2 showcases the capability of RGPSSM for hyperparameter and inducing point adaptation in a synthetic wing rock dynamics learning task. Section 6.3 evaluates the learning performance of RGPSSM on various real-world datasets.
The learning algorithm is implemented on a desktop computer running Python 3.9, with an Intel(R) Core(TM) i7-14700F 2.10 GHz processor and 32 GB of RAM, and the associated code is published online [^8] https://github.com/TengjieZheng/rgpssm. In all experiments, the multi-output kernel (33) is employed, where the single-output kernel is the Squared Exponential Automatic Relevance Determination (SEARD) [15] kernel, which is defined as:
@@
where the diagonal matrix $\bm{\Lambda}=\mathrm{diag}(l_{1}^{2},l_{2}^{2},\dots,l_{d_{x}}^{2})$ encodes the length scales for each input dimension. The signal covariance matrix $\bm{\Xi}$ is defined as $\mathrm{diag}(\sigma_{1}^{2},\sigma_{2}^{2},\dots,\sigma_{d_{f}}^{2})$, where each $\sigma_{i}^{2}$ denotes the signal variance corresponding to the $i$-th output dimension of the function. This kernel formulation enables the model to flexibly characterize both the relevance of different input features (via $l_{i}$) and the signal strength of each output (via $\sigma_{i}^{2}$). All these hyperparameters will be optimized online to adaptively capture the underlying function characteristics during learning.
@@ 6.1 Synthetic NASCAR® Dynamics Learning
To benchmark the online learning performance of RGPSSM, we compare it with two state-of-the-art algorithms: SVMC [44], which represents PF-based methods, and OEnVI [39], which exemplifies SVI-based methods. Since the pre-parameterized AKF-based method is a special case of the proposed method, it is not included in the comparison. In the experiment, all methods are tested using synthetic NASCAR® data [60], which involves a two-dimensional state with dynamics following a recurrent switching linear dynamical system. The measurement model is given by $\bm{y}_{t}=\bm{C}\bm{x}_{t}+\bm{w}_{t},\,\bm{w}_{t}\sim N(\bm{0},0.1^{2}\bm{I}_{4})$, where $d_{y}=4$ and $\bm{C}$ is a 4-by-2 matrix with random elements. The three methods are trained with 500 measurements and tested by predicting the subsequent 500 steps. Note that, unlike the experiments in [44] and [39], we use fewer dimensions and a smaller number of measurements for training, providing a more rigorous evaluation. For a fair comparison, the maximum number of inducing points across all methods is limited to 20, and the GP hyperparameters are set to the same values and are not adjusted online. The implementation and other parameter settings for the SVMC and OEnVI algorithms are based on the code provided online [^9] https://github.com/catniplab/svmc  [^10] https://github.com/zhidilin/gpssmProj.
@@ 6.2 Synthetic Wing Rock Dynamics Learning
where the system state is defined as $\bm{x}_{t}=[\theta(t),\,p(t)]^{\top}$, with $\bm{x}_{t}^{(2)}=p(t)$, the control input $c_{t}=\delta_{a}(t)$, and the uncertainty is modeled as $f(\bm{x}_{t})=\Delta(\theta(t),p(t))$. To enable hyperparameter adaptation, we perform one iteration of hyperparameter optimization during each RGPSSM update using a learning rate of 0.01. The number of inducing points is limited to 20. Under these settings, we evaluate the proposed method and analyze the impact of hyperparameter and inducing point adaptation.
Firstly, we conduct a Monte Carlo test of the proposed learning algorithm with different initial values of GP hyperparameters. The simulation results are depicted in Fig. 3 and 4. As shown in Fig. 3, all three GP hyperparameters gradually converge near the reference values, which are obtained through offline GP training using data pairs $(\theta,p,\Delta)$. This indicates that the GP hyperparameters do not need to be precisely selected at the stage of algorithm initialization, and the online optimization can effectively adjust them. To evaluate the improvement resulting from hyperparameter adaptation, we calculate the prediction RMSE over the last quarter segment, which is 29.3% lower with hyperparameter adaptation than without. Furthermore, to more intuitively illustrate the benefits of hyperparameter adaptation, we present the state profile from one Monte Carlo simulation case, as shown in Fig. 4. In the figure, the blue solid lines depict the true values of state $\theta(t),p(t)$ and uncertainty $\Delta(t)$; the orange dashed lines represent the filtered state and uncertainty predictions obtained without hyperparameter adaptation; and the green dashed lines show the predictions with hyperparameter adaptation. It is evident that hyperparameter adaptation improves both the mean and variance predictions of the GP, thereby enhancing the filtering accuracy of the state, especially the roll rate $p$. Therefore, the proposed method can not only learn the uncertain dynamics but also improve the state estimation accuracy.
@@ 7 Conclusion
In this paper, we propose a novel recursive learning method, RGPSSM, for online inference in Gaussian Process State Space Models (GPSSMs). First, to enable learning in any operating domain, we derive a two-step Bayesian update equation that does not rely on any GP approximation, which provides a new solution for the nonlinearity and nonparametric issues of GPSSMs. Second, to mitigate the computational burden imposed by the nonparametric nature of GPs, we introduce a dynamic inducing-point set adjustment algorithm to ensure computational efficiency. Third, to eliminate the need for careful tuning of GP hyperparameters at initialization and to enhance online learning accuracy, we develop an online optimization method that extracts measurement information from the approximate posterior distribution. Fourth, we derive a Cholesky version of RGPSSM in detail to further improve numerical stability, and also present the extension of RGPSSM to the multi-output GP case. Fifth, we analyze the computational complexity of the proposed method and demonstrate that it generalizes several existing approaches, as well as highlight its strong potential for further extensions. Finally, through extensive evaluations on a diverse set of real and synthetic datasets, we demonstrate the superior learning performance and adaptability of the proposed method. In the future, research may focus on enhancing nonlinear filtering accuracy, improving computational efficiency, scaling it to high-dimensional problems, and adapting it to process and measurement noise covariances, as well as time-varying dynamics.
Appendix
7.1 Score for Selection of Inducing Point

### v4

@@ 4.3 Algorithm Summary
In summary, we have addressed the four challenges outlined in the Introduction for implementing online GPSSMs. Specifically, first, nonlinearity and coupling are handled through first-order linearization and joint inference of the system state and GP, leading to a closed-form Bayesian update equation. Second, a dynamic inducing-point adjustment algorithm is developed to overcome the computational burden caused by the non-parametric nature of the GP. Third, we recover the likelihood model from the approximate posterior distribution and utilize it for the online optimization of GP hyperparameters, thus enhancing learning accuracy. Based on these solutions, we have developed an efficient online GPSSM method that can adapt to both the operating domain and GP hyperparameters. Compared to existing online GPSSM methods, the proposed method is the first to achieve recursive inference for GPSSMs with the ability to adapt to both the operating domain and GP hyperparameters. This makes it truly applicable to various flexible learning scenarios. Therefore, we refer to this method as the recursive GPSSM (RGPSSM). The pseudocode of RGPSSM is summarized in Algorithm 1, and the schematic diagram of the algorithm workflow is shown in Fig. 1.
@@
1: Input: initial inducing points $\bm{u}$, threshold for adding points $\varepsilon_{tol}$, budget for the inducing-point set $M$.
2: Repeat until operation ends:
3:   Assess the novelty $\gamma$ of the new point $f_{t}$ using (23).
4:   If $\gamma>\varepsilon_{tol}$ then
5:     Add $f_{t}$ to the inducing-point set and propagate the moments using (17).
6:   Else
7:     Propagate the moments using (24).
@@
10:   If a new measurement $y_{t}$ is available then
11:     Correct the moments using (18).
12:    Minimize $\mathcal{L}$ in (32) to update the GP hyperparameters $\bm{\theta}$ using the Adam optimizer [41], and subsequently update the moments using (30).
@@
In theory, Algorithm 1 has recursive learning capability for GPSSMs. However, when implemented on a computer, there exists numerical stability issues as in the Kalman filter. Specifically, the algorithm requires the evolution of the covariance matrix $\bm{\Sigma}_{t}$ of the augmented state $\bm{X}_{t}$, which can easily lose positive definiteness due to the accumulation of rounding errors. Compared to general KF applications, this issue is relatively severe in the context of RGPSSM because the dimension of the joint covariance $\bm{\Sigma}_{t}$ is often higher. To address this problem, a stable implementation based on Cholesky factorization is derived, as detailed in Appendix D. This implementation method is capable of enhancing the numerical stability by making the computation more compact. In the next section, we will discuss the computational complexity of the Cholesky version of RGPSSM, along with the extension to the multi-output GP cases.
@@ 5.1 Multi-output GP Version
where ${K}_{0}(\cdot,\cdot)$ in the RHS represents the single-output kernel, and $\bm{\Xi}$ denotes the signal covariance between each function output dimension. According to the multi-output GP literature [42], this kernel belongs to the linear model of coregionalization (LMC) method. Based on this multi-output kernel function, the kernel matrix has a Kronecker product structure, i.e., $\bm{K}_{0}(\mathcal{X}_{1},\mathcal{X}_{2})={K}_{0}(\mathcal{X}_{1},\mathcal{X}_{2})\otimes\bm{\Xi}$. This structure indicates that the $(i,j)$-th $d_{f}\times d_{f}$ block of the kernel matrix $\bm{K}_{0}(\mathcal{X}_{1},\mathcal{X}_{2})$ is the prior covariance between function values $\bm{f}(\textbf{x}_{i})$ and $\bm{f}(\textbf{x}_{j})$. Therefore, the only modification from the single-output GP version of RGPSSM to the multi-output GP version is to convert scalar elements in related vectors or matrices into corresponding blocks, which are presented in the following.
@@
  • $\bm{\Omega}_{dd}$ is the $i_{\mathrm{d}}$-th diagonal block of the inducing-point submatrix in $\Omega$, where the inducing-point submatrix is defined as the matrix slice $\bm{\Omega}[d_{x}:,d_{x}:]$
@@
Therefore, the extension to the multi-output GP version is straightforward, with modifications involving the kernel function, the definitions of some vectors and matrices, and the evaluation criteria for inducing points management. It is evident that the dimensions of the statistical moments scale with the output dimension of the GP, thus impacting the computational complexity. This is analyzed in the next subsection.
@@ 5.2 Computational Complexity Analysis
The computational complexity of RGPSSM is analyzed for its various algorithmic steps in Table 1, which presents a detailed comparison between the standard and Cholesky-based implementations of RGPSSM. This analysis is based on the multi-output GP version of RGPSSM.
@@
Based on the complexity analysis in Table 1, the following additional remarks are provided:
  • By precomputing and storing the Cholesky factor of the kernel matrix $\bm{K}_{uu}$ during updates, we can avoid repeated matrix inversions in GP prediction (10), thereby reducing the computational complexity of the GP prediction to quadratic order.
@@
  • It is clear that the overall algorithmic complexity grows cubically with both the function output dimension $d_{f}$ and the number of inducing points $M$. This is mainly due to two factors: the EKF-like moment matching, which provides second-order learning speed [43], and the flexible inducing point adjustment and hyperparameter optimization algorithms. Although this cubic complexity limits the algorithm’s applicability to high-dimensional scenarios, it offers a stable, fast, and flexible learning solution for low-dimensional problems, which are common in many applications such as aerospace [44] and robotics [36].
@@ 5.3 Generalization and Applicability
RGPSSM can be regarded as a unified generalization of several existing methods. Specifically, it subsumes both the pre-parameterized AKF-based online GPSSM with inducing points approximation [28] and kernel-based online regression approaches such as KRLS [37] and SOGP [39] as special cases. For example, if RGPSSM is initialized with a fixed set of inducing points and no further updates are performed (i.e., by setting $\varepsilon_{\text{tol}}=+\infty$), it reduces to the pre-parameterized AKF-based method. Furthermore, when the transition and measurement functions in the GPSSM are set as $x_{t}=f(\bm{c}_{t})$ and $y_{t}=x_{t}$, where $\bm{c}_{t}$ and $y_{t}$ denote the function input and output, respectively, RGPSSM becomes equivalent to KRLS and SOGP. Therefore, RGPSSM offers practical and flexible recursive learning capabilities applicable to a wide range of scenarios, as summarized below:
  • Applicable to various observation scenarios, including partial state observation ($\bm{y}_{t}=\bm{g}(\bm{x}_{t})$), full state observation ($\bm{y}_{t}=\bm{x}_{t}$), and direct function value observation ($\bm{y}_{t}=\bm{f}(\bm{x}_{t})$), as well as any combination of these cases.
@@
  • The use of first-order linearization, together with the augmentation of the inducing-point set, enables a finite-dimensional Gaussian approximation of the joint state–inducing point distribution. This provides a flexible foundation for incorporating advanced nonlinear filters, such as the sigma point filter [45, 46], iterative filter [47], variational filter [48], or exact moment evaluation [49], to further improve the accuracy of moment matching.
@@ 6 Experiment Result
This section demonstrates the proposed method using two synthetic simulations and a real-world dataset. Section 6.1 presents a comparative experiment on a synthetic NASCAR® dataset, which serves as a benchmark for online GPSSM methods. Section 6.2 showcases the capability of RGPSSM for hyperparameter and inducing point adaptation in a synthetic wing rock dynamics learning task. Section 6.3 evaluates the learning performance of RGPSSM on real-world system identification tasks.
The learning algorithm is implemented on a desktop computer running Python 3.9, with an Intel(R) Core(TM) i7-14700F 2.10 GHz processor and 32 GB of RAM, and the associated code is published online [^3] https://github.com/TengjieZheng/rgpssm. In all experiments, the multi-output kernel (33) is employed, where the single-output kernel is the commonly used Squared Exponential Automatic Relevance Determination (SEARD) [14] kernel, which is defined as:
@@
where the diagonal matrix $\bm{\Lambda}=\mathrm{diag}(l_{1}^{2},l_{2}^{2},\dots,l_{d_{x}}^{2})$ encodes the length scales for each input dimension. The signal covariance matrix $\bm{\Xi}$ is defined as $\mathrm{diag}(\sigma_{1}^{2},\sigma_{2}^{2},\dots,\sigma_{d_{f}}^{2})$, where each $\sigma_{i}^{2}$ denotes the signal variance corresponding to the $i$-th output dimension of the function. This kernel formulation enables the model to flexibly characterize both the input-output relevance of different features (via $l_{i}$) and the signal strength of each output (via $\sigma_{i}^{2}$). All these hyperparameters will be optimized online to adaptively capture the underlying function characteristics during learning.
@@ 6.1 Synthetic NASCAR® Dynamics Learning
To benchmark the online learning performance of RGPSSM, we compare it with two state-of-the-art (SOTA) online GPSSM algorithms: SVMC [30], which represents the PF-based methods, and OEnVI [26], which exemplifies the SVI-based methods. Since the pre-parameterized AKF-based method is a special case of the proposed method, it is not included in the comparison. In the experiment, all methods are tested using synthetic NASCAR® data [50]. This dataset is a benchmark for online GPSSM methods and is used in [30] and [26]. This dataset involves a two-dimensional state with dynamics that follow a recurrent switching linear dynamical system (see [50] for more details). The transition model is $\bm{x}_{t+1}=\bm{x}_{t}+\bm{f}(\bm{x}_{t})$, where the function $\bm{f}(\bm{x}_{t})$ is defined by the dataset in [50] and is learned by the GP. The measurement model is given by $\bm{y}_{t}=\bm{C}\bm{x}_{t}+\bm{w}_{t}$, where $\bm{w}_{t}\sim N(\bm{0},0.1^{2}\bm{I}_{4})$, $d_{y}=4$, and $\bm{C}$ is a 4-by-2 matrix with random elements. The three methods are trained with 500 measurements and tested by predicting the subsequent 500 steps. Note that, unlike the experiments in [30] and [26], we use fewer dimensions and a smaller number of measurements for training, providing a more rigorous evaluation. For a fair comparison, the maximum number of inducing points across all methods is limited to 20, and the GP hyperparameters are set to the same values and are not adjusted online. The implementation and other parameter settings for the SVMC and OEnVI algorithms are based on the code provided online [^4] https://github.com/catniplab/svmc [^5] https://github.com/zhidilin/gpssmProj.
@@ 6.2 Synthetic Wing Rock Dynamics Learning
where the system state is defined as $\bm{x}_{t}=[\theta(t),\,p(t)]^{\top}$, with $\bm{x}_{t}^{(2)}=p(t)$, the control input $c_{t}=\delta_{a}(t)$, and the uncertainty is modeled as $f(\bm{x}_{t})=\Delta(\theta(t),p(t))$. To enable hyperparameter adaptation, we perform one iteration of hyperparameter optimization during each RGPSSM update using a learning rate of 0.01. The number of inducing points is limited to 20. Under these settings, we evaluate the proposed method and analyze the impact of adapting to the hyperparameter and inducing point.
Firstly, we conduct a Monte Carlo test of the proposed learning algorithm with different initial values of GP hyperparameters. The simulation results are depicted in Fig. 3 and 4. As shown in Fig. 3, all three GP hyperparameters gradually converge near the reference values, which are obtained through offline GP training using data pairs $(\theta,p,\Delta)$. This indicates that the GP hyperparameters do not need to be precisely selected at the stage of algorithm initialization, and the online optimization can effectively adjust them. To evaluate the improvement resulting from hyperparameter adaptation, we calculate the prediction RMSE over the last quarter segment, which is 29.3% lower with hyperparameter adaptation than without. Furthermore, to more intuitively illustrate the benefits of hyperparameter adaptation, we present the state profile from one Monte Carlo simulation case, as shown in Fig. 4. It is evident that hyperparameter adaptation improves both the mean and variance predictions of the GP, thereby enhancing the filtering accuracy of the state, especially the roll rate $p$. Therefore, the proposed method can not only learn the uncertain dynamics but also improve the state estimation accuracy.
@@ 7 Conclusion
In this paper, we propose a novel recursive learning method, RGPSSM, for online inference in Gaussian Process State Space Models (GPSSMs). First, to enable learning in any operating domain, we derive a two-step Bayesian update equation that does not rely on any GP approximation, which provides a new solution for the nonlinearity and nonparametric issues of GPSSMs. Second, to mitigate the computational burden imposed by the nonparametric nature of GPs, we introduce a dynamic inducing-point set adjustment algorithm to ensure computational efficiency. Third, to eliminate the need for careful tuning of GP hyperparameters at initialization and to enhance online learning accuracy, we develop an online optimization method that extracts measurement information from the approximate posterior distribution. Fourth, we derive a Cholesky version of RGPSSM in detail to further improve numerical stability, and also present the extension of RGPSSM to the multi-output GP case. Fifth, we analyze the computational complexity of the proposed method and demonstrate that it generalizes several existing approaches, as well as highlight its strong potential for further extensions. Finally, through extensive evaluations on a diverse set of real and synthetic datasets, we demonstrate the superior learning performance and adaptability of the proposed method.
8 Acknowledgement
This work is supported by the National Natural Science Foundation of China (Grant No. 12572052) and the Beijing Natural Science Foundation (Grant L251013).
Appendix A Score for Selection of Inducing Point